# Universal Video Temporal Grounding with Generative Multi-modal Large Language Models

### v1

@@ 2.2 Universal Temporal Grounding Model
where $\phi_{\text{vision}}(\cdot)$ is the vision encoder, which supports dynamic input resolutions with a fixed patch size of $P$, and $\phi_{\text{project}}(\cdot)$ is the projection module. $\psi_{\text{resize}}(\cdot)$ denotes the frame resizing function, which adjusts the spatial resolution of each frame so that it can be partitioned into $N_{\text{res}}$ patches of size $P\times P$ pixels. $\psi_{\text{compress}}(\cdot)$ denotes token compression via bilinear interpolation, yielding $N_{\text{res}}$ tokens as the final representation. Unlike frame resizing, which may degrade spatial details, token compression reduces redundancy at the token level while better preserving key semantic information, especially for long videos with low resolutions per frame. When the number of frames exceeds $N_{f}^{\text{long}}$, the video is divided into multiple clips of length $N_{f}^{\text{long}}$ for divide-and-conquer processing.
@@
Timestamp-Interleaved Sequence Construction. To associate frames with their timestamps, we encode temporal information by interleaving timestamp tokens between visual tokens. For a frame $f_{i}$ sampled at timestamp $t_{i}$, the timestamp is represented as free text $\tau_{i}=$ ‘‘${}t_{i}$ seconds’’, and inserted before visual tokens of $f_{i}$, forming an interleaved sequence:
@@
where $\phi_{\text{tokenizer}}(\cdot)$ is the tokenizer. This interleaved sequence is input to the large language models (LLMs), which generates temporal boundaries for a given query $Q$, with the output formatted as $\mathcal{A}=$‘‘From $s_{k}$ seconds to $e_{k}$ seconds’’. Note that, unlike conventional methods that predict the exact moment in the original video, our model instead predicts the minimal timestamp interval $(s_{k},e_{k})$ from the sampled timestamps set $\mathcal{T}$. That is to say, we leverage the retrieval capabilities of MLLMs to read out the inserted timestamp tokens, rather than decoding dense position encodings.
@@
Discussion. Here, explicitly providing timestamps offers several crucial advantages: (i) Improved temporal grounding. These explicit timestamp tokens help the LLM effectively associate visual content with its corresponding temporal position in the video, allowing it to generalize to temporal grounding tasks. (ii) Alignment-free integration. As the inserted timestamps are textual, they naturally reside in the language space and do not require realignment. In contrast, methods based on learned vectors necessitate additional embedding alignment, thereby complicating the LLM’s understanding of temporal information. (iii) Generalisable temporal reasoning. When combined with adaptive frame scaling, explicitly provided timestamp tokens allow the model to directly refer to the temporal cues, potentially enabling extrapolation. For example, as demonstrated in Section 3.5, this design allows the model to be trained on short or medium-length videos (e.g., around 6 minutes), while still generalizing effectively to much longer videos (e.g., 20 minutes) during inference on downstream tasks. (iv) Model-agnostic design. This method is plug-and-play and can be easily incorporated into diverse model architectures with minimal modifications, eliminating the need for architecture-specific adjustments such as custom positional encoding schemes guo2025vtg ; bai2025qwen2 .
@@
Building on these designs, our framework achieves universal temporal grounding by processing input frames at different spatial resolutions and generating temporal predictions at appropriate granularities, according to video duration. For long videos, we employ a hierarchical, multi-stage inference strategy. Specifically, if a video exceeds $N_{f}^{\text{long}}$ frames, it is partitioned into multiple clips, each of length $N_{f}^{\text{long}}$. Segment retrieval is first performed within each clip to identify potentially relevant segments. The retrieved segments from all clips are then aggregated and subjected to further segment retrieval, and this process can be repeated recursively if necessary. Finally, fine-grained grounding is conducted within the final selected segments. As evidenced by Sec 3.3, this flexible and hierarchical inference process ensures accurate and efficient temporal grounding across diverse video durations.
@@ 3.1 Experimental Setups
3.2 Comparison with State-of-the-arts on Video Temporal Grounding
@@ 3.4 Robustness Test
UniVTG 9.80 6.20 7.76 50.00 1.37 4.92 3.86 34.87 62.56 54.67
VTimeLLM 3.87 1.40 3.68 48.60 1.62 1.22 1.58 34.60 59.52 54.30
TimeSuite 6.28 0.20 1.22 48.80 5.06 2.89 2.09 32.47 58.51 53.25
@@ 5 Conclusion
In the appendix, we provide additional experimental details (Appendix A), implementation details (Appendix B), qualitative results (Appendix C), more ablation studies (Appendix D), discussions of limitations and future work (Appendix E), and broader impacts (Appendix F).
@@ C.2 Qualitative Results for ANet-Captions
Appendix D More Ablation Studies
D.1 Effect of Model Size.
To assess the impact of MLLM capacity, we ablate model size on Ego4D-NLQ and Charades-STA. As Table 8 shows, the larger Qwen2-VL-7B outperforms its smaller counterpart Qwen2-VL-2B. This suggests that increased model capacity enhances multimodal reasoning, thereby further improving temporal grounding performance when integrated with our proposed method.
Table 8: Ablation on model size. All experiments are conducted under the dataset-specific setting.
MLLMs Size Charades-STA Ego4D-NLQ
R1@.5 R1@.7 mIoU R1@.3 R1@.5 mIoU
Qwen2-VL 2B 65.38 42.18 57.25 10.50 5.80 7.29
Qwen2-VL 7B 74.33 53.71 63.15 24.79 16.83 17.25
D.2 Effect of Model Type.
As discussed in Section 2.2, our approach is compatible with any MLLM that can process dynamic frame inputs. To assess architectural impact, we evaluate two MLLM variants: Qwen2-VL-7B and Qwen2.5-VL-7B. As shown in Table 9, our method consistently achieves superior temporal grounding performance regardless of the underlying MLLM. These results demonstrate that the effectiveness of our framework can continually benefit from advances in MLLM architectures, enabling further improvements in temporal grounding as more powerful MLLMs become available.
Table 9: Ablation on model type. All experiments are conducted under the dataset-specific setting.
MLLMs Charades-STA ANet-Captions QVHighlights Ego4D-NLQ TaCoS
R1@.5 R1@.7 R1@.5 R1@.7 R1@.5 R1@.7 R1@.3 R1@.5 R1@.3 R1@.5
Qwen2-VL-7B 74.33 53.71 54.81 36.62 77.76 63.29 24.79 16.83 61.18 48.31
Qwen2.5-VL-7B 74.19 53.98 55.64 37.64 77.17 63.10 23.53 16.12 61.61 49.16
Appendix E Limitations & Future Work
@@
Appendix F Broader Impacts Statement

### v2

@@ 2.2 Universal Temporal Grounding Model
where $\phi_{\text{vision}}(\cdot)$ is the vision encoder, which supports dynamic input resolutions with a fixed patch size of $P$, and $\phi_{\text{project}}(\cdot)$ is the projection module. $\psi_{\text{resize}}(\cdot)$ denotes the frame resizing function, which adjusts the spatial resolution of each frame so that it can be partitioned into $N_{\text{res}}$ patches of size $P\times P$ pixels. $\psi_{\text{compress}}(\cdot)$ denotes token compression via bilinear interpolation, yielding $N_{\text{res}}$ tokens as the final representation. Unlike frame resizing, which may degrade spatial details, token compression reduces redundancy at the token level while better preserving key semantic information, especially for long videos with low resolutions per frame. When the number of frames exceeds $N_{f}^{\text{long}}$, the video is divided into multiple clips of length $N_{f}^{\text{long}}$ for divide-and-conquer processing. The ablation experiments of token budget, thresholds and the input processing strategy can be found in Appendix E.3 and E.1.
@@
[…]
@@
where $\phi_{\text{tokenizer}}(\cdot)$ is the tokenizer. This interleaved sequence is input to the large language models (LLMs), which generates temporal boundaries for a given query $Q$, with the output formatted as $\mathcal{A}=$‘‘From $s_{k}$ seconds to $e_{k}$ seconds’’. Note that, unlike conventional methods that predict the exact moment in the original video, our model instead predicts the minimal timestamp interval $(s_{k},e_{k})$ from the sampled timestamps set $\mathcal{T}$. That is to say, we leverage the retrieval capabilities of MLLMs to read out the inserted timestamp tokens, rather than decoding dense position encodings. For a comparison of these temporal information encoding approaches, refer to the Appendix E.2.
@@
Discussion. Here, explicitly providing timestamps offers several crucial advantages: (i) Improved temporal grounding. These explicit timestamp tokens help the LLM effectively associate visual content with its corresponding temporal position in the video, allowing it to generalize to temporal grounding tasks. (ii) Alignment-free integration. As the inserted timestamps are textual, they naturally reside in the language space and do not require realignment. In contrast, methods based on learned vectors necessitate additional embedding alignment, thereby complicating the LLM’s understanding of temporal information. (iii) Generalisable temporal reasoning. When combined with adaptive frame scaling, explicitly provided timestamp tokens allow the model to directly refer to the temporal cues, potentially enabling extrapolation. For example, as demonstrated in Section 3.7, this design allows the model to be trained on short or medium-length videos (e.g., around 6 minutes), while still generalizing effectively to much longer videos (e.g., 20 minutes) during inference on downstream tasks. (iv) Model-agnostic design. This method is plug-and-play and can be easily incorporated into diverse model architectures with minimal modifications, which is validated in Section 3.4, eliminating the need for architecture-specific adjustments such as custom positional encoding schemes guo2025vtg; bai2025qwen2.
@@
Building on these designs, our framework achieves universal temporal grounding by processing input frames at different spatial resolutions and generating temporal predictions at appropriate granularities, according to video duration. For long videos, we employ a hierarchical, multi-stage inference strategy. Specifically, if a video exceeds $N_{f}^{\text{long}}$ frames, it is partitioned into multiple clips, each of length $N_{f}^{\text{long}}$. Segment retrieval is first performed within each clip to identify potentially relevant segments. The retrieved segments from all clips are then aggregated and subjected to further segment retrieval, and this process can be repeated recursively if necessary. Finally, fine-grained grounding is conducted within the final selected segments. As evidenced by Section 3.5, this flexible and hierarchical inference process ensures accurate and efficient temporal grounding across diverse video durations.
@@ 3.1 Experimental Setups
3.2 Comparison with State-of-the-arts on Video Temporal Grounding
@@
Table 5: Closed-source Model Evaluation. All models are evaluated on sampled subsets, with outputs lacking predicted timestamps excluded.
Model Ego4D-NLQ Charades-STA
R1@.3 R1@.5 mIoU R1@.5 R1@.7 mIoU
Gemini-2.5-flash comanici2025gemini 15.29 15.29 13.93 36.46 13.54 38.72
Gemini-2.5-pro comanici2025gemini 26.83 19.51 20.45 35.96 11.24 39.28
GPT-4.1-mini openai2025gpt4.1 5.00 3.00 5.04 31.00 17.00 31.46
GPT-4o openai2024gpt4o 10.59 7.06 7.97 41.11 17.78 43.82
Seed1.5-VL seed2025seed1_5vl 19.39 10.20 13.74 90.82 63.27 73.69
UniTime-Full 25.00 17.00 17.20 81.00 61.00 70.11
Table 6: Flexibility Verification. Improvements over baseline are highlighted in (green). All UniTime variants use the data-specific setting.
Method Ego4D-NLQ Charades-STA
R1@.3 R1@.5 mIoU R1@.5 R1@.7 mIoU
Qwen2-VL-2B wang2024qwen2 0.34 0.34 0.42 2.23 0.51 6.70
      +UniTime 10.50 5.80 7.29 (+6.87) 65.38 42.18 57.25 (+50.55)
Qwen2-VL-7B wang2024qwen2 0.48 0.17 0.46 7.18 3.28 8.98
      +UniTime 24.79 16.83 17.25 (+16.79) 74.33 53.71 63.15 (+54.17)
Qwen2.5-VL-7B bai2025qwen2 1.11 0.48 0.82 60.32 34.27 52.42
      +UniTime 23.53 16.12 16.61 (+15.79) 74.19 53.98 63.62 (+11.20)
InternVL2.5-2B chen2024expanding 0.33 0.22 0.42 5.48 1.77 9.30
      +UniTime 9.87 5.26 7.46 (+7.04) 69.49 47.77 59.03 (+49.73
InternVL2.5-8B chen2024expanding 2.19 0.55 1.77 8.66 2.55 17.54
      +UniTime 16.23 8.55 12.03 (+10.26) 71.83 49.62 60.50 (+42.96)
3.3 Comparison with Closed-source models on Video Temporal Grounding
To comprehensively assess UniTime, we benchmarked mainstream closed-source MLLMs to examine whether UniTime surpasses them or narrows the gap on VTG tasks. We conducted temporal grounding experiments on a randomly sampled subset from Ego4D and Charades. As shown in Table 5, Seed1.5-VL seed2025seed1_5vl attained the strongest performance on short videos (72.2 mIoU on Charades), while Gemini-2.5-Pro comanici2025gemini exhibited superior grounding on long videos (20.5 mIoU on Ego4D). However, these closed-source models struggle to achieve high performance concurrently on both long and short videos, indicating a lack of generalization. In contrast, UniTime achieves competitive performance compared to these closed-source models while maintaining more balanced capabilities for both short and long-form video grounding benchmarks, proving our approach to be overall the best in terms of consistent performance across varying video types.
3.4 Flexibility Verification
As discussed in Section 2.2, our approach is compatible with any MLLM that processes dynamic frame inputs. To evaluate flexibility and generalization, we conduct experiments across multiple baseline MLLMs (e.g., Qwen2-VL, Qwen2.5-VL, InternVL2.5) and model scales. As shown in Table 6, larger MLLMs outperform their smaller counterparts, and our method consistently improves temporal grounding performance regardless of the underlying architecture. These results demonstrate strong generalization and flexibility, and indicate that the effectiveness of our framework continues to benefit from advances in MLLM architectures and increased model scale, enabling further gains in temporal grounding as more powerful models become available.
@@ 3.6 Robustness Test
Recently, a related work jung2025consistency focuses on prediction consistency: they also curate test data by rephrasing original queries and shifting ground-truth moments in videos to probe grounding consistency. For detailed tests and comparisons, please refer to the Appendix D.
@@
UniVTG lin2023univtg 9.80 6.20 7.76 50.00 1.37 4.92 3.86 34.87 62.56 54.67
VTimeLLM huang2024vtimellm 3.87 1.40 3.68 48.60 1.62 1.22 1.58 34.60 59.52 54.30
TimeSuite zeng2024timesuite 6.28 0.20 1.22 48.80 5.06 2.89 2.09 32.47 58.51 53.25
@@ 5 Conclusion
Acknowledgements
This work is supported by National Key R&\&D Program of China (No.2022ZD0161400).
@@ Acknowledgements
In the appendix, we provide additional experimental details (Appendix A), implementation details (Appendix B), qualitative results (Appendix C), grounding consistency evaluation (Appendix D), more ablation studies (Appendix E), discussions of limitations and future work (Appendix F), and broader impacts (Appendix G).
@@ C.2 Qualitative Results for ANet-Captions
Appendix D Grounding Consistency Evaluation
To further evaluate UniTime’s robustness in temporal grounding, we conduct grounding consistency assessment on Charades-CON and ActivityNet-CON jung2025consistency, which are curated from subsets of CharadesSTA and ActivityNet-Captions, respectively. The grounding consistency targets two complementary consistency dimensions:
[…]
ii) Shifted Grounding (S-Ground). This metric assesses whether a model consistently grounds the same visual content when that content is shifted to different temporal positions. Specifically, while preserving the internal frame sequence of the ground-truth moment $m_{\mathrm{gt}}$, it is randomly shifted to a new moment $m_{s}$, with evaluations conducted on the resulting predictions.
Experimental results, summarized in the Table 10, demonstrate that UniTime not only exhibits strong temporal grounding performance (Ground) but also achieves excellent consistency and robustness (R-Ground and S-Ground). UniTime attains performance comparable to VTune jung2025consistency, which extends instruction tuning with explicit consistency objectives, and in some cases surpasses it.
Table 10: Grounding consistency evaluation of Video-MLLMs. Relative consistency scores are in brackets.
Method Charades-CON ActivityNet-CON
Ground R-Ground S-Ground Ground R-Ground S-Ground
GPT-4o openai2024gpt4o 28.5 21.2 (74.3) 9.3 (32.8) 26.8 18.1 (67.5) 10.4 (38.8)
Gemini 1.5 Flash team2024gemini 34.6 29.7 (85.7) 24.8 (71.7) 37.8 30.8 (81.4) 24.8 (65.6)
Video-LLaMA zhang2023video 14.2 10.6 (74.9) 5.3 (37.6) 12.8 8.5 (66.8) 7.2 (56.8)
Video-LLaMA + VTune jung2025consistency 54.4 38.2 (70.3) 10.9 (20.0) 33.0 24.7 (74.8) 10.0 (30.2)
TimeChat ren2024timechat 30.5 25.0 (82.1) 5.6 (18.5) 4.6 2.9 (64.1) 1.0 (21.2)
TimeChat + VTune jung2025consistency 76.2 69.2 (90.8) 36.2 (47.5) 37.4 28.3 (75.6) 10.6 (28.3)
UniTime-Full 81.5 73.9 (90.7) 58.3 (71.6) 49.1 30.2 (61.5) 21.3 (43.4)
Appendix E More Ablation Studies
E.1 Effect of Video Processing Strategies
To justify our design choice in Section 2.2, i.e., resizing short videos and compressing long videos, we conduct an ablation study on the Ego4D-NLQ dataset. This study evaluates how different video processing strategies impact both grounding performance and computational efficiency.
Table 11: Ablation on video processing strategies. Seg Retrieval Acc. indicates the accuracy of coarse-grained segment retireval.
Short Video Long Video R1@.3 R1@.5 mIoU Seg Retrieval Acc. Training Time
Frame Resizing Token Compression 24.79 16.83 17.25 45.28 21 h
Frame Resizing Frame Resizing 18.53 12.67 12.88 35.36 12 h
Token Compression Token Compression 24.06 16.18 16.71 45.28 24.5 h
Our findings indicate that: (i) For long videos, adaptive frame scaling reduces tokens per frame while keeping the frame rate fixed. Though some visual cues are lost, this feature-level token compression retains more semantic information than frame resizing, which discards input resolution. This leads to better segment retrieval accuracy and recall (row 1 vs. row 2). (ii) For short clips with sufficient tokens, both methods perform similarly (row 1 vs. row 3). However, token compression requires high-resolution inputs before merging, incurring higher computational costs. In UniTime-Full, using token compression for short videos increased training time by 5 days compared to frame resizing.
Therefore, rather than employing a unified strategy, we adopt a hybrid approach to balance performance and efficiency: for short videos, we apply frame resizing to reduce computational overhead, while for long videos, we leverage token compression to better preserve visual details. This design choice enables our method to achieve strong performance with practical training costs.
E.2 Comparison of Temporal Information Encoding
We conduct ablation experiments on Charades-STA to compare the effectiveness of Timestamp Token Insertion versus Dense Position Encodings.
Table 12: Ablation on temporal information encoding.
Method R1@.5 R1@.7 mIoU
Dense Position Encodings 48.44 27.15 44.85
Timestamp Token Insertion (UniTime-SP) 74.33 53.71 63.15
Our base model, Qwen2-VL-7B, utilizes Multimodal Rotary Position Embedding (MRoPE), which decomposes position embeddings into temporal, height, and width components. After fine-tuning, temporal grounding remains suboptimal with dense position encodings. In contrast, timestamp token insertion yields significant gains. We attribute this improvement to the MLLM’s strong multimodal retrieval ability and the simplification of the task into (i) identifying question-relevant video segments and (ii) directly reading their corresponding timestamp tokens.
E.3 Ablation Study of Adaptive Frame Scaling Hyperparameters
In this section, we conduct an ablation study on the threshold hyperparameters $N_{f}^{\text{long}}$, $N_{f}^{\text{short}}$, and the token budget $N_{\text{total}}$ within Adaptive Frame Scaling, elucidating the rationale behind our parameter choices.
Threshold $N_{f}^{long}$. This parameter specifies the maximum video length that can be processed in a single pass. For videos with length $\leq N_{f}^{\text{long}}$, we process the input directly; for videos with length $N_{f}^{\text{long}}$, we partition the input into segments, process each segment independently, and then aggregate the predictions. As shown in Table 13, varying $N_{f}^{\text{long}}$ has only a marginal effect on accuracy because over-length inputs are segmented and subsequently fused. Moreover, Adaptive Frame Scaling maintains a fixed token budget per segment, so reducing $N_{f}^{\text{long}}$ increases the number of segments and the total token count at inference, thereby incurring higher latency. Balancing efficiency and accuracy, we set $N_{f}^{\text{long}}=1024$.
Threshold $N_{f}^{short}$. This hyperparameter determines whether inference is executed in a single stage or via a multi-stage (coarse-to-fine) pipeline. For inputs with temporal length $\leq N_{f}^{\text{short}}$, we use single-stage processing; for inputs with temporal length $N_{f}^{\text{short}}$, we employ multi-stage processing. As shown in Table 14, within the 64–128 s duration range, the performance difference between multi-stage ($N_{f}^{\text{short}}{=}64$) and single-stage ($N_{f}^{\text{short}}\geq 128$) configurations is modest. However, a lower threshold increases average inference latency by routing more inputs to the multi-stage pipeline. For inputs in the 128–256 s range, multi-stage processing yields substantially higher accuracy (54.57 vs. 42.90 mIoU). To balance accuracy and efficiency, we set $N_{f}^{\text{short}}{=}128$.
Token budget $N_{\text{total}}$. This parameter specifies the maximum number of tokens permitted. As shown in Table 15, performance improves monotonically as the token budget increases, up to the available GPU capacity. Accordingly, we set $N_{\text{total}}=16,384$, which corresponds to the upper limit supported by our hardware.
Table 13: Ablation of the threshold $N_{f}^{\text{long}}$ on Ego4D.
$N_{f}^{long}$ R1@.3 R1@.5 mIoU
1024 24.79 16.83 17.25
512 24.72 16.63 16.95
256 23.94 16.10 16.68
Table 14: Ablation of the threshold $N_{f}^{\text{short}}$ on TaCoS.
$N_{f}^{short}$ mIoU mIoU mIoU
(64-128s) (128-256s) Avg.
64 58.43 54.57 50.28
128 56.59 54.57 50.00
256 56.59 42.90 46.82
Table 15: Ablation of the token budget $N_{\text{total}}$ on Ego4D.
Budget R1@.3 R1@.5 mIoU
16384 24.79 16.83 17.25
8192 24.08 16.10 16.69
4096 23.04 15.07 15.96
Appendix F Limitations & Future Work
@@
Appendix G Broader Impacts Statement